# OneDiff: A Generalist Model for Image Difference Captioning

### v3

@@ 3.1 Task-Specific Data Collection
We firstly curate data from existing datasets that focus on identifying differences between image pairs, as is shown in Fig. 2. These include benchmarks like Spot-the-Diff [12] and CLEVR-Change, which are directly compatible with our IDC tasks. Additionally, the Birds-to-Words dataset, which is tailored for describing differences between birds, and the Image-Editing-Request dataset, which provides instructions for editing images, are also included. The NLVR2 dataset, known for its dual-image inputs, further enhances the model’s ability to capture intricate details in image pairs.
@@ 5.3 Performance Comparison
As shown in Tab. 1(a), we evaluate the performance of our model on Image-Editing-Request. OneDiff demonstrates significant superiority in this benchmark. Specifically, compared to the SOTA models, the improvement of our OneDiff on CIDEr is nearly doubled, i.e. 211% compare with SCORER+CBR. We then compare the performance on Spot-the-Diff. As detailed in Tab. 1(b), OneDiff achieves significant improvements across BLEU, CIDEr and METEOR compared with SOTAs. Finally, we analyze the performance disparities on the Birds-to-Words. Consistent with the previous two benchmarks, our model demonstrated a significant lead as shown in Tab. 1(c). Quantitative results comparing to other IDC methods are displayed in in Fig. 5. OneDiff generates more reasonable difference captions with correct details on the bad cases of these methods.
@@ 5.4 Ablation Analysis
We first assess the efficacy of the proposed CST strategy and VDM module. We report the related results in Tab. 2 with the 6-layer and 32-delta-token VDM. The hyper-parameter setting of VDM is discussed in our supplementary materials. Specifically, on one hand, the CST strategy facilitates cost-effective and efficient adaptation for model to multi-image understanding domains, yielding improvements across all benchmarks. On the other hand, while the isolated implementation of the VDM module does not significantly outperform the baseline (which we attribute to insufficient training of the VDM due to the CST being disabled during the pretrain phase), the concurrent activation of both CST and VDM substantially boosts performance demonstrating the synergistic benefits, significantly surpassing the performance of a single strategy. We opt for a more efficient language model in Tab. 3(a), reducing computational costs while maintaining commendable performance on Birds-to-Words. To further demonstrate the role of VDT, we randomly drop a portion of them during inference to test the impact on performance on Birds-to-Words in Tab. 3(b).
@@
In this section, we systematically analyze the impact of both realistic and synthetic data generated via our proposed pipeline, as indicated by the performance metrics outlined in Tab. 4. Specifically, incorporating either type of data independently yield consistent performance gains across three benchmarks, which attests to the high quality and extensive distribution of our data. Moreover, utilizing both data types in tandem further enhance performance, underscoring the complementary nature of the two data patterns. Overall, the data we constructed have substantially augmented the existing IDC dataset, effectively mitigating the data scarcity issue in this field.
@@ Appendix 0.A Appendix
  • Sec. 0.A.1 Ablation experiments onthe hyper-parameters settings of the VDM module in our OneDiff framework.
  • Sec. 0.A.2 The statistics of our proposed DiffCap Dataset, both real-image and synthetic-image samples from DiffCap and some potential failure cases.
@@ 0.A.1 Ablation
The hyper-parameter setting of VDM is explored in Tab. 5. For efficiency, we only train 3000 steps in Stage II when exploring the model design. The number of delta tokens varies in 16, 32, 64 and 128 with 6 or 12 Transformer layers. OneDiff is not quite sensitive to different hyper-parameters with relatively stable performances.
@@ 0.A.2 Data
Our DiffCap dataset consists of three components: collected datasets, generated data from real images and generated data from synthetic images. (See Tab. 6)
@@
More samples from DiffCap are displayed as qualitative results of date construction. (See Fig. 6 and Fig. 7)
@@ 0.A.2.3 Potential Failure Cases
We present potential failure cases where the environment and objectives are roughly the same, yet multiple describable image differences exist, which could lead to confusion. (See Fig. 8)

### v4

@@ 3.1 Task-Specific Data Collection
We firstly curate data from existing datasets that focus on identifying differences between image pairs, as is shown in 2. These include benchmarks like Spot-the-Diff [12] and CLEVR-Change, which are directly compatible with our IDC tasks. Additionally, the Birds-to-Words dataset, which is tailored for describing differences between birds, and the Image-Editing-Request dataset, which provides instructions for editing images, are also included. The NLVR2 dataset, known for its dual-image inputs, further enhances the model’s ability to capture intricate details in image pairs.
@@ 5.3 Performance Comparison
As shown in 1(a), we evaluate the performance of our model on Image-Editing-Request. OneDiff demonstrates significant superiority in this benchmark. Specifically, compared to the SOTA models, the improvement of our OneDiff on CIDEr is nearly doubled, i.e. 211% compare with SCORER+CBR. We then compare the performance on Spot-the-Diff. As detailed in 1(b), OneDiff achieves significant improvements across BLEU, CIDEr and METEOR compared with SOTAs. Finally, we analyze the performance disparities on the Birds-to-Words. Consistent with the previous two benchmarks, our model demonstrated a significant lead as shown in 1(c). Quantitative results comparing to other IDC methods are displayed in in 5. OneDiff generates more reasonable difference captions with correct details on the bad cases of these methods.
@@ 5.4 Ablation Analysis
We first assess the efficacy of the proposed CST strategy and VDM module. We report the related results in 2 with the 6-layer and 32-delta-token VDM. The hyper-parameter setting of VDM is discussed in our supplementary materials. Specifically, on one hand, the CST strategy facilitates cost-effective and efficient adaptation for model to multi-image understanding domains, yielding improvements across all benchmarks. On the other hand, while the isolated implementation of the VDM module does not significantly outperform the baseline (which we attribute to insufficient training of the VDM due to the CST being disabled during the pretrain phase), the concurrent activation of both CST and VDM substantially boosts performance demonstrating the synergistic benefits, significantly surpassing the performance of a single strategy. We opt for a more efficient language model in 3(a), reducing computational costs while maintaining commendable performance on Birds-to-Words. To further demonstrate the role of VDT, we randomly drop a portion of them during inference to test the impact on performance on Birds-to-Words in 3(b).
@@
In this section, we systematically analyze the impact of both realistic and synthetic data generated via our proposed pipeline, as indicated by the performance metrics outlined in 4. Specifically, incorporating either type of data independently yield consistent performance gains across three benchmarks, which attests to the high quality and extensive distribution of our data. Moreover, utilizing both data types in tandem further enhance performance, underscoring the complementary nature of the two data patterns. Overall, the data we constructed have substantially augmented the existing IDC dataset, effectively mitigating the data scarcity issue in this field.
@@ Appendix 0.A Appendix
  • 0.A.1 Ablation experiments onthe hyper-parameters settings of the VDM module in our OneDiff framework.
  • 0.A.2 The statistics of our proposed DiffCap Dataset, both real-image and synthetic-image samples from DiffCap and some potential failure cases.
@@ 0.A.1 Ablation
The hyper-parameter setting of VDM is explored in 5. For efficiency, we only train 3000 steps in Stage II when exploring the model design. The number of delta tokens varies in 16, 32, 64 and 128 with 6 or 12 Transformer layers. OneDiff is not quite sensitive to different hyper-parameters with relatively stable performances.
@@ 0.A.2 Data
Our DiffCap dataset consists of three components: collected datasets, generated data from real images and generated data from synthetic images. (See 6)
@@
More samples from DiffCap are displayed as qualitative results of date construction. (See 6 and 7)
@@ 0.A.2.3 Potential Failure Cases
We present potential failure cases where the environment and objectives are roughly the same, yet multiple describable image differences exist, which could lead to confusion. (See 8)